# Exact Work Distribution and Jarzynski’s Equality of a Relativistic Particle in an Expanding Piston

### v1

@@ Abstract
We study the non-equilibrium work in a pedagogical model of relativistic ideal gas. We obtain the exact work distribution and verify the Jarzynski’s equality. In the non-relativistic limit, our results recover the non-relativistic results [Lua et al, J. Phys. Chem. B, 109, 6805(2005)]. We also find that, unlike the non-relativistic case, the work distribution no longer has zeros and the number of collisions in this relativistic gas model is finite. In addition, based on an analysis of the experimental parameters, we conclude that it is difficult to detect the relativistic effects of the work distribution of the ideal gas in a piston system with the current experimental techniques.
@@ I Introduction
The Jarzynski’s equality (JE) [1] is one of the most elegant results in the field of non-equilibrium statistical physics. While being a direct result from Liouvelle’s theorem and supposed to hold in arbitrary Hamiltonian systems including relativistic systems, the work distribution of a relativistic system under an arbitrary work protocol has not been explored so far. In this article, we study the work distribution of a simple relativistic model consisting of only a one-dimensional cylinder, a piston and relativistic ideal gas. Previous work on this model has been done and the JE verified, but in a very limited range of parameters [2] such that the collision only happens no more than once for all particles. Moreover, we still lack the detailed information of work distribution of the system giving rise to various fluctuation theorems, which plays a fundamental role in non-equilibrium statistical physics. A Newtonian approximation of our model [3], however, has yielded analytic results both for the JE and the work distribution, suggesting the solvability in the relativistic regime. Relevant to the non-relativistic ideal gas in a piston, the non-equilibrium work distribution of quantum gas in an expanding piston [4][5] has been studied and the JE verified. It is desirable to extend those studies to the relativistic regime. In this article, we analytically compute the work distribution and verify the JE in this simple setup. Our model can be viewed as a relativistic generalization of the one in Ref. [3] and recovers it in the low-speed and low-temperature limit. Such a toy model would be helpful for understanding more complicated and more realistic systems’ behavior under relativistic conditions. As we will see in this paper, although the work distribution changes drastically in the extreme relativistic regime, the equality itself is independent of the microscopic dynamics. In the following, we will study the non-equilibrium work distribution of relativistic gas in an expanding piston. The results can serve as a pedagogical example and provide intuitive insights to the robustness of the JE. Also, we will show that it is difficult to detect the relativistic effect of the work distribution with the current experimental techniques.
@@
The article is organized as follows: In Sec. II, we demonstrate the validity of the JE under the framework of relativistic mechanics. In Sec. III, we analytically calculate the work distribution of a relativistic particle in an expanding piston system. In the non-relativistic limit, we recover the results in Ref. [3]. In Sec. IV, we give the discussion and summary.
II Jarzynski’s equality under the framework of relativistic mechanics
The original proof of the JE was based mainly on the classical Hamilton mechanics and its corollary, i. e., Liouville’s theorem [1]. In this section, we extend the proof to the framework of relativistic mechanics. We take the spacetime dimension to be $1+1$ to suit our model. Note that the generalization to higher spacetime dimensions is straightforward. The manifestly covariant dynamics for a particle of the static mass $m$ can be formulated as
$f^{\mu}=m\frac{\mathrm{d}u^{\mu}}{\mathrm{d}\zeta},$ (1)
$-\frac{\partial V}{\partial q}=\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{mu}{\sqrt{1-u^{2}/c^{2}}}\right),$ (2)
where $V$ denotes the potential, $c$ the speed of light, and $q,u,t$ are the position, velocity and time measured in the laboratory frame. With this form of dynamics, we can easily construct the Hamiltonian of a system of $N$ particles of mass $m_{1},\,m_{2},\,...,\,m_{N}$. Furthermore, we let the Hamiltonian be controlled by an external agent via the parameter $\lambda=\lambda(t)$ of the potential $V_{\lambda}.$ The Hamiltonian reads
where $q_{i}$ is the position of the $i$-th particle and $p_{i}$ its conjugate momentum.
The canonical equations are
As long as the canonical equations are formulated, one can easily generalize Liouville’s theorem to relativistic regime [6]. The theorem states that the Jacobian determinant of the canonical coordinates $(p_{i,t},q_{i,t})$ at time $t$ as functions of the initial canonical coordinates $(p_{i,0},q_{i,0})$ at time 00 is
The initial equilibrium state of the inverse temperature $\beta$ is determined by a probability distribution $\rho(p_{i,0},q_{i,0})=\exp{[-\beta H(p_{i,0},q_{i,0})]}/Z_{0}$ with $Z_{0}=\int\mathrm{d}p_{i,0}\mathrm{d}q_{i,0}\exp[-\beta H_{\lambda(0)}(p_{i,0},q_{i,0})]$ being the initial partition function. We will see in the following (Subsec. III.2) that for ideal gas the distribution is the so-called Maxwell-Jüttner distribution [7]. The proof of the JE follows by calculating the expectation value of the exponential work $\exp{[W(p_{i},q_{i},\tau)]}$ done by the system along the trajectory up to time $\tau.$ By definition we have
where the Liouville’s theorem is used for the third equality and $Z_{\tau}$ denotes the final equilibrium state partition function. The partition functions can be expressed in the form of free energy (see, for example, [8]), resulting in
$\left\langle e^{\beta W}\right\rangle=e^{-\beta(F_{\tau}-F_{0})},$ (9)
where $F_{\tau}$ and $F_{0}$ are the free energy of the equilibrium states corresponding to $\lambda(\tau)$ and $\lambda(0)$. Thus, we demonstrate the validity of the JE under the framework of relativistic mechanics.
III A Relativistic Piston Model
@@
The model we consider here is nothing more than some ideal gas inside a one-dimensional cylinder. Suppose initially the length of the vessel is $L$, and the gas is of the inverse temperature $\beta$. We now expand the piston outwards at the speed $v_{p}$, and stop it after a time interval $\tau$.
@@
where $W_{1}$ and $\Delta F_{1}$ denote the work and the change of the free energy per particle. We can now see that for the ideal gas, the single-particle quantities $W_{1}$ and $\Delta F_{1}$ also satisfy the JE. Thus, we limit our ideal gas to just a single particle, and from now on we omit the subscript $1$ for the discussion of the single-particle quantities.
III.1 Trajectory of a Single Particle
To calculate the trajectory work as a function of the initial state, one must at first figure out how the trajectory of a single particle is like.
Let us denote the velocity of the particle after the $n$-th collision with the moving piston as $v_{n}.$ Following the moving of a particle, we find out that after the $(n+1)$-th collision with the moving piston, the speed of the particle is reduced to
@@ III.1 Trajectory of a Single Particle
This result can be derived simply by performing Lorentz transformation twice, noting that the piston is an inertial reference frame during the whole process and the collisions are elastic collisions. The solution to the recurrence relation of the particle’s speed $v_{n}$ after the $n$-th collision can also be derived
$v_{n}=\frac{(c+v)\alpha_{p}^{2n}-c+v}{(c+v)\alpha_{p}^{2n}+c-v}\cdot c,$ (12)
@@
Another thing we must figure out is the time $t_{n}$ when the $n$-th collision with the moving piston takes place. We have the recurrence relation
@@
Here $x$ denotes the initial position of the particle. Note that the sign of $x$ can be either positive or negative due to the initial velocity can be either towards or away from the moving piston. So from now on we simply extend the range of $x$ from $[0,L]$ to $[-L,L]$ to remove the negative sign (for details, please see Appendix A). With the expression of $v_{n}$, the product of a sequence can be simplified
@@
$\begin{split}t_{n}=&\left[(-\alpha_{p}^{2n}-\alpha_{p}+\alpha_{p}^{n+1}+\alpha_{p}^{n})\frac{v}{c}\right.\\ &\left.+(-\alpha_{p}^{n+1}+\alpha_{p}^{n})\frac{x}{L}+(\alpha_{p}-\alpha_{p}^{2n})\right]\\ &\left[-\alpha_{p}+\alpha_{p}^{2n}+(\alpha_{p}+\alpha_{p}^{2n})\frac{v}{c}\right]^{-1}\cdot\frac{L(1+\alpha_{p})}{c(1-\alpha_{p})},\end{split}$ (17)
which is the time of the $n$-th collision between the particle and the moving piston. It is obvious that $n$ can not take an arbitrarily large number, and the $n$-th collision is guaranteed if and only if both
@@
The first requirement ensures that the collision happens before the ending time $\tau$, and the second that the particle can catch up with the moving piston. These requirements give a maximum number of collisions
@@
where []delimited-[][$\cdots$][ ⋯ ] denotes the integer part of $\cdots$.⋯ .
@@
III.2 A Direct Verification of Jarzynski’s Equality in the Expanding Relativistic Piston Model
It is worth emphasizing that the model we considered here is not characterized by a time-dependent Hamiltonian, but by a parameterized boundary condition [9]. Thus the proof of the JE in Sec. II is inapplicable to the expanding rigid piston system. Still, we can demonstrate that the JE is valid in the expanding rigid piston system. In order to do so, we focus on the Jacobian determinant between the initial and the final state and show it to be unity.
@@ III.2 A Direct Verification of Jarzynski’s Equality in the Expanding Relativistic Piston Model
where $K_{1}$ is the modified Bessel function of the second kind, $(x,p)$ are the initial position and the initial momentum, and $k_{B}$ is the Boltzmann’s constant. This distribution can also be expressed in the position-velocity space, with the initial velocity denoted by $v$, as
@@
at time $\tau.$ The Jaccobian determinant can be directly computed from Eq. (24),
@@
III.3 Distribution of Work
There are three dimensionless parameters in our model: $k_{B}T/(mc^{2}),$ $L/(c\tau)$ and $v_{p}/c.$ Therefore it is convenient to set $m=c=k_{B}=L=1,$ leaving only $v_{p},$ $\tau$ and $\beta=1/T$ as free parameters.
Using the probability distribution Eq. (21), we can evaluate the distribution of work $W$
@@ III.3 Distribution of Work
After some tedious calculations (see Appendix B for details), the distribution function of $W$ can be analytically expressed as
@@
$T_{n}=\frac{\alpha_{p}^{-(n-1)}-1-\alpha_{p}+\alpha_{p}^{n}}{1-\alpha_{p}}+\frac{\alpha_{p}^{-(n-1)}+\alpha_{p}^{n}}{1+\alpha_{p}}\cdot\tau,$ (31)
$X_{n}=\frac{\alpha_{p}^{-(n-1)}-\alpha_{p}^{n}}{1-\alpha_{p}}+\frac{\alpha_{p}^{-(n-1)}-\alpha_{p}^{n}}{1+\alpha_{p}}\cdot\tau.$ (32)
Here, $\xi_{n}(v)$ is a function of $v$ and denotes the initial position of the particles that happen to collide with the moving piston exactly $n$ times within time $\tau$ with the initial velocity $v$. For a pictorial explanation, please see Appendix B. Note that $\xi_{n}(v)$ is a linear function of $v$. This is not as obvious as in the non-relativistic regime. For an intuitive explanation, please see Appendix A.
@@
III.4 Non-relativistic Limit
In the non-relativistic limit, we expect our results to recover those in Ref. [3]. For our choice of units, where the speed of light is set to be 1, the Newtonian mechanics is recovered by taking $v_{p}\ll 1$ and $W\sim 1/\beta\sim v_{p}^{2}.$
We will deal with the exponential part in Eq. (27) separately. For now we can expand the rest part of the work distribution except the exponential and the normalization constant at $v_{p}$ around 00 and $W$ around $0$ We have
$\begin{split}&[1-\xi_{n}(v_{n}(W))]\cdot\frac{1}{(\alpha_{p}^{-n}-1)\left[1+\alpha_{p}^{n}-v_{n}(W)(1-\alpha_{p}^{n})\right]}\\ &=-\frac{n-1}{2nv_{p}}-\frac{(n-1)\tau}{4n}+\frac{\tau W}{8n^{2}v_{p}^{2}}+O(v_{p}^{2}),\end{split}$ (34)
$2\cdot\frac{1}{(\alpha_{p}^{-n}-1)\left[1+\alpha_{p}^{n}-v_{n}(W)(1-\alpha_{p}^{n})\right]}=\frac{1}{2nv_{p}},$ (35)
@@ III.4 Non-relativistic Limit
$\begin{split}&[1+\xi_{n+1}(v_{n}(W))]\cdot\frac{1}{(\alpha_{p}^{-n}-1)\left[1+\alpha_{p}^{n}-v_{n}(W)(1-\alpha_{p}^{n})\right]}\\ &=\frac{n+1}{2nv_{p}}+\frac{(n+1)\tau}{4n}-\frac{\tau W}{8n^{2}v_{p}^{2}}+O(v_{p}^{2})\end{split}$ (36)
@@
The exponential term left in Eq. (27), together with the normalization constant $K_{1}(\beta)$, needs to be treated with extra care. We start by noting that the thermodynamic temperature $\beta$ is of the order $v_{p}^{-2},$ so
$-\frac{\beta}{\sqrt{1-v_{n}(W)^{2}}}=-\beta-\frac{\beta}{2}\left(\frac{W}{2nv_{p}}+nv_{p}\right)^{2}+O(v_{p}),$ (37)
where a temperature dependent constant $e^{-\beta}$ appears. The low-temperature limit also affects the constant factors, yielding
$\frac{e^{-\beta}}{K_{1}(\beta)}\approx\sqrt{\frac{2\beta}{\pi}}$ (38)
as $\beta$ approaches $\infty$, which is exactly the Maxwellian normalization constant.
To conclude, we are able to recover the non-relativistic results in Ref. [3] ( $\tau=1$ is chosen in accordance with Ref. [3] )
@@
In Fig. 0(a)-0(c) we plot the work distributions of the expanding relativistic piston model and its non-relativistic limit. The deviations of the relativistic results from the non-relativistic ones with different choices of parameters are shown. One can see that as expected, the relativistic results deviate more prominently from the non-relativistic results at high temperature and fast speed.
(a) $\tau=10,\,v_{p}=0.1,\,\beta=3$
(b) $\tau=100,\,v_{p}=0.01,\,\beta=10$
(c) $\tau=1000,\,v_{p}=0.001,\,\beta=33$
Figure 1: Relativisitic and non-relativistic work distribution with different parameters. Take hydrogen atoms as an example. Supposing that the length of the cylinder is $L=1\mathrm{cm}$ (which is much larger than the thermal length of the atoms). The parameters are chosen to be $\tau=0.3\mathrm{ns},v_{p}=3\times 10^{7}\mathrm{m/s},T=3\times 10^{12}\mathrm{K}$ in (a), $\tau=3\mathrm{ns},v_{p}=3\times 10^{6}\mathrm{m/s},T=1\times 10^{12}\mathrm{K}$ in (b), and $\tau=30\mathrm{ns},v_{p}=3\times 10^{5}\mathrm{m/s},T=3\times 10^{11}\mathrm{K}$ in (c).
III.5 Limit of Fast-moving Piston
We have already known that in Newtonian mechanics, at a very large $v_{p}$, the validity of the JE relies on the far tails of the Maxwellian distribution [3]. It is thus intriguing to also think of this problem in special relativity, where every speed has the speed of light $c$ as its upper bound. The main obstacle to the application of the JE is that to measure the average of exponential work, one must repeat the experiment a certain number of times. The exponent makes sure that the contribution of the tail of the distribution, while its probability goes to zero, is non-vanishing [10]. Specifically in our model, when the speed of the moving piston approaches the speed of light, the fraction of particles that can collide with the piston approaches 0. In such a case an experiment with non-zero work is of probability

### v2

@@ Abstract
We study the nonequilibrium work in a pedagogical model of relativistic ideal gas. We obtain the exact work distribution and verify Jarzynski’s equality. In the nonrelativistic limit, our results recover the nonrelativistic results of Lua and Grosberg [J. Phys. Chem. B, 109, 6805(2005)]. We also find that, unlike the nonrelativistic case, the work distribution no longer has zeros and the number of collisions in this relativistic gas model is finite. In addition, based on an analysis of the experimental parameters, we conclude that it is difficult to detect the relativistic effects of the work distribution of the ideal gas in a piston system with the current experimental techniques.
@@ I Introduction
Jarzynski’s equality (JE) [1] is one of the most elegant results in the field of nonequilibrium statistical physics. While it is a direct result from Liouville’s theorem and is supposed to hold in arbitrary Hamiltonian systems including relativistic systems, the work distribution of a relativistic system under an arbitrary work protocol has not been explored so far. In this article we study the work distribution of a simple relativistic model consisting of only a one-dimensional cylinder, a piston, and relativistic ideal gas. Previous work on this model has been done and the JE verified, but in a very limited range of parameters [2] such that the collision only happens no more than once for all particles. Moreover, we still lack detailed information of the work distribution of the system giving rise to various fluctuation theorems, which plays a fundamental role in nonequilibrium statistical physics. A Newtonian approximation of our model [3], however, has yielded analytic results for both the JE and the work distribution, suggesting the solvability in the relativistic regime. Relevant to the nonrelativistic ideal gas in a piston, the nonequilibrium work distribution of quantum gas in an expanding piston [4, 5] has been studied and the JE verified. It is desirable to extend those studies to the relativistic regime. In this article, we analytically compute the work distribution and verify the JE in this simple setup. Our model can be viewed as a relativistic generalization of the one in Ref. [3] and recovers it in the low-speed and low-temperature limit. Such a toy model would be helpful for understanding more complicated and more realistic systems’ behavior under relativistic conditions. As we will see in this paper, although the work distribution changes drastically in the extreme relativistic regime, the equality itself is independent of the microscopic dynamics. In the following, we will study the nonequilibrium work distribution of relativistic gas in an expanding piston. The results can serve as a pedagogical example and provide intuitive insights into the robustness of the JE. Also, we will show that it is difficult to detect the relativistic effect of the work distribution with the current experimental techniques.
@@
The article is organized as follows:. In Sec. II we analytically calculate the work distribution of a relativistic particle in an expanding piston system and verify the JE. In the nonrelativistic limit, we recover the results in Ref. [3]. In Sec. III we discuss the main results and summarize our work.
II Relativistic Piston Model
@@
The model we consider here is nothing more than some ideal gas consisting of $N$ molecules inside a one-dimensional cylinder. Suppose initially the length of the vessel is $L$ and the gas is initially at the inverse temperature $\beta$, where $\beta=1/k_{B}T$, with $k_{B}$ the Boltzmann constant and $T$ the temperature. We now expand the piston outward at the speed $v_{p}$ and stop it after a time interval $\tau$.
@@
where $W_{1}$ and $\Delta F_{1}$ denote the work and the change of the free energy per particle, respectively. We can now see that for the ideal gas, the single-particle quantities $W_{1}$ and $\Delta F_{1}$ also satisfy the JE. Thus, we limit our ideal gas to just a single particle, and from now on we omit the subscript $1$ for the discussion of the single-particle quantities.
II.1 Trajectory of a single particle
To calculate the trajectory work as a function of the initial state, we must first determine the trajectory of a single particle.
Let us denote the velocity of the particle after the $n$th collision with the moving piston by $v_{n}.$ Following the moving of a particle, we find out that after the $(n+1)$th collision with the moving piston, the speed of the particle is reduced to
@@ II.1 Trajectory of a single particle
This result can be derived simply by performing Lorentz transformation twice, noting that the piston is an inertial reference frame during the whole process and the collisions are elastic collisions. When $v_{n},v_{p}\ll c,v_{n+1}=\left(v_{n}-2v_{P}+v_{n}v_{p}^{2}/c^{2}\right)/\left(1+v_{p}^{2}/c^{2}-2v_{n}v_{p}/c^{2}\right)\approx v_{n}-2v_{p}$, which is the result in the nonrelativistic limit.
The solution to the recurrence relation of the particle’s speed $v_{n}$ after the $n$th collision can also be derived as
$v_{n}=\frac{(c+v)\alpha_{p}^{2n}-c+v}{(c+v)\alpha_{p}^{2n}+c-v}c,$ (3)
@@
Another thing we must figure out is the time $t_{n}$ when the $n$th collision with the moving piston takes place. We have the recurrence relation
@@
Here $x$ denotes the initial position of the particle. The sign of $x$ is positive if the initial velocity is towards the moving piston and negative if the initial velocity is away from the moving piston. So from now on we simply extend the range of $x$ from $[0,L]$ to [L,L][-L,L] to remove the negative sign (for details, see Appendix A). With the expression of $v_{n}$, the product of a sequence can be simplified as
@@
$\begin{split}t_{n}=&\left[(-\alpha_{p}^{2n}-\alpha_{p}+\alpha_{p}^{n+1}+\alpha_{p}^{n})\frac{v}{c}\right.\\ &\left.+(-\alpha_{p}^{n+1}+\alpha_{p}^{n})\frac{x}{L}+(\alpha_{p}-\alpha_{p}^{2n})\right]\\ &\times\left[-\alpha_{p}+\alpha_{p}^{2n}+(\alpha_{p}+\alpha_{p}^{2n})\frac{v}{c}\right]^{-1}\cdot\frac{L(1+\alpha_{p})}{c(1-\alpha_{p})},\end{split}$ (8)
which is the time of the $n$th collision between the particle and the moving piston.
The speed of the particle cannot be larger than $c$. Therefore, during a finite period $\tau$, $n$ cannot take an arbitrarily large number, and the $n$th collision is guaranteed if and only if both
@@
The first requirement ensures that the collision happens before the ending time $\tau$ and the second that the particle can catch up with the moving piston. The particles with the initial velocity $v\to c$ and initial position $x=L$ undergo the maximum number of collisions. With these conditions and letting $t_{n}\leq\tau$ in Eq. (8), we find the maximum number of collisions
@@
where $\cdots$] denotes the integer part of $\cdots.$
@@
II.2 Direct verification of Jarzynski’s equality in the expanding relativistic piston model
Jarzynski’s equality, as a result of Liouville’s theorem, should be directly generalized to any Hamiltonian systems. In Appendix B we give a proof of the JE in a relativistic system. It is worth emphasizing that the model we consider here is not characterized by a time-dependent Hamiltonian, but by a parametrized boundary condition [6]. Thus the proof of the JE in Appendix B is inapplicable to the expanding rigid piston system. Still, we can demonstrate that the JE is valid in the expanding rigid piston system. In order to do so, we focus on the Jacobian determinant between the initial and the final state and show it to be unity.
@@ II.2 Direct verification of Jarzynski’s equality in the expanding relativistic piston model
where $K_{1}$ is the modified Bessel function of the second kind, $(x,p)$ are the initial position and the initial momentum, and $m$ is the static mass of the particle. This distribution can also be expressed in the position-velocity space, with the initial velocity denoted by $v$, as
@@
at time $\tau.$ With Eq. (3), (8), and (14), the Jacobian determinant can be directly computed from Eq. (15) (for details, see Appendix C),
@@
II.3 Distribution of work
There are three dimensionless parameters in our model: $\beta mc^{2},$ $L/c\tau$ and $v_{p}/c.$ Therefore it is convenient to set $m=c=k_{B}=L=1,$ leaving only $v_{p},$ $\tau$ and $\beta$ as free parameters.
Using the probability distribution (12), we can evaluate the distribution of work $W$
@@ II.3 Distribution of work
After some tedious calculations (see Appendix D for details), the distribution function of $W$ can be analytically expressed as
@@
$T_{n}=\frac{\alpha_{p}^{-(n-1)}-1-\alpha_{p}+\alpha_{p}^{n}}{1-\alpha_{p}}+\frac{\alpha_{p}^{-(n-1)}+\alpha_{p}^{n}}{1+\alpha_{p}}\tau,$ (22)
$X_{n}=\frac{\alpha_{p}^{-(n-1)}-\alpha_{p}^{n}}{1-\alpha_{p}}+\frac{\alpha_{p}^{-(n-1)}-\alpha_{p}^{n}}{1+\alpha_{p}}\tau.$ (23)
Here, $\xi_{n}(v)$ is a function of $v$ and denotes the initial position of the particles that happen to collide with the moving piston exactly $n$ times within time $\tau$ with the initial velocity $v$. For a pictorial explanation, see Appendix D. Note that $\xi_{n}(v)$ is a linear function of $v$. This is not as obvious as in the nonrelativistic regime. For an intuitive explanation, see Appendix A.
@@
II.4 Nonrelativistic Limit
In the nonrelativistic limit, we expect our results to recover those in Ref. [3]. Here it is convenient to restore the speed of light $c$, and the nonrelativistic limit thus corresponds to $c$ approaching infinity.
To restore $c$ one needs to do some dimensional analysis. With the presence of $c$, the dimensions of the relevant physical quantities are $W]=\mathbf{T}^{-2}$, $v_{n}(W)]=\mathbf{T}^{-1}$, $P(W)]=\mathbf{T}^{2}$, $\beta]=\mathbf{T}^{2}$, $\tau]=\mathbf{T}$, and $c]=\mathbf{T}^{-1}$, where $\mathbf{T}$ denotes the dimension of time. We see that to restore $c$ one only needs to make the substitutions $W\mapsto W/c^{2},$ $v_{n}(W)\mapsto v_{n}(W)/c,$ $\beta\mapsto\beta c^{2},$ and $P(W)\mapsto P(W)c^{2}.$ We then have
$\begin{split}P(W)=&\frac{P_{0}\delta(W/c^{2})}{c^{2}}+\frac{1}{2K_{1}(\beta c^{2})c^{2}}\sum_{n=1}^{N}\varphi_{n}\left(\frac{v_{n}(W/c^{2})}{c}\right)\\ \times&\frac{e^{-(\beta c^{2})/\sqrt{1-v_{n}(W/c^{2})^{2}/c^{2}}}}{(\alpha_{p}^{-n}-1)\left[1+\alpha_{p}^{n}-v_{n}(W/c^{2})(1-\alpha_{p}^{n})/c\right]}.\end{split}$ (25)
We will deal with the exponential part in Eq. (18) separately. For now we can expand the rest part of the work distribution except the exponential and the normalization constant at $c$ around infinity. We have
[…]
$\begin{split}&\frac{2}{(\alpha_{p}^{-n}-1)\left[1+\alpha_{p}^{n}-v_{n}(W/c^{2})(1-\alpha_{p}^{n})/c\right]}\\ &=\frac{c}{2nv_{p}}+O(1),\end{split}$ (27)
@@ II.4 Nonrelativistic Limit
$\begin{split}&\frac{1+\xi_{n+1}\left(v_{n}(W/c^{2})/c\right)}{(\alpha_{p}^{-n}-1)\left[1+\alpha_{p}^{n}-v_{n}(W/c^{2})(1-\alpha_{p}^{n})/c\right]}\\ &=\frac{(n+1)c}{2nv_{p}}+\frac{(n+1)c\tau}{4n}-\frac{c\tau W}{8n^{2}v_{p}^{2}}+O(1)\end{split}$ (28)
@@
The exponential term in Eq. (18), together with the normalization constant $K_{1}(\beta)$, needs to be treated with extra care. We start by taking the limit $c\to\infty$ of the exponential
$\begin{split}&-\frac{\beta c^{2}}{\sqrt{1-v_{n}\left(W/c^{2}\right)^{2}/c^{2}}}\\ &=-\beta c^{2}-\frac{\beta}{2}\left(\frac{W}{2nv_{p}}+nv_{p}\right)^{2}+O(c^{-2}),\end{split}$ (29)
where a temperature dependent constant $e^{-\beta c^{2}}$ appears as the leading term. The limit of $c\to\infty$ also affects the constant factors, yielding
$\frac{e^{-\beta c^{2}}}{K_{1}(\beta c^{2})}\approx\sqrt{\frac{2\beta}{\pi}}c,$ (30)
which is exactly the Maxwellian normalization constant multiplied by $c$. Multiplying all pieces together we see that the $c$ dependence cancels, which agrees with the physical intuition that the low-speed physics is independent of the speed of light.
To conclude, we are able to recover the nonrelativistic results in Ref. [3] ($\tau=1$ is chosen in accordance with Ref. [3])
@@
In Fig. 1(a)-3(a) we plot the work distributions of the expanding relativistic piston model and its nonrelativistic limit. The deviations of the relativistic results from the nonrelativistic ones with different choices of parameters are shown. One can see that as expected, the relativistic results deviate more prominently from the nonrelativistic results at high temperature and fast speed.
Figure 3: Relativistic and nonrelativistic work distribution with different parameters. We use hydrogen atoms as an example and suppose that the length of the cylinder is $L=1~\mathrm{cm}$ (which is much larger than the thermal length of the atoms). The parameters are chosen to be (a) $\tau=0.3~\mathrm{ns},v_{p}=3\times 10^{7}~\mathrm{m/s},T=3\times 10^{12}~\mathrm{K}$, and $\beta=3$; (b) $\tau=3~\mathrm{ns},v_{p}=3\times 10^{6}~\mathrm{m/s},T=1\times 10^{12}~\mathrm{K}$, and $\beta=10$; and (c) $\tau=30~\mathrm{ns},v_{p}=3\times 10^{5}~\mathrm{m/s},T=3\times 10^{11}~\mathrm{K}$, and $\beta=33$
II.5 Limit of sast-moving piston
We already know that in Newtonian mechanics, at a very large $v_{p}$, the validity of the JE relies on the far tails of the Maxwellian distribution [3]. It is thus intriguing to also think of this problem in special relativity, where every speed has the speed of light $c$ as its upper bound. The main obstacle to the application of the JE is that to measure the average of exponential work, one must repeat the experiment a certain number of times. The exponent makes sure that the contribution of the tail of the distribution, while its probability goes to zero, is nonvanishing [8]. Specifically in our model, when the speed of the moving piston approaches the speed of light, the fraction of particles that can collide with the piston approaches 0. In such a case an experiment with nonzero work is of probability